OpenStudy (anonymous):

Complete the square. 2x^(2)+6x+y^(2)+10y+100=0

OpenStudy (anonymous):

\[\implies2x^2 + 6x + y^2 + 10y + 100 \]\[\implies (2x^2 + 6x) + (y^2 + 10y) = -100\]\[\implies 2(x^2 + 3x) + (y^2 + 10y) = -100 \]\[\implies 2(x^2 + 3x + 2.25) + (y^2 + 10y + 25) = -100 + 4.5 + 25\]\[\implies 2(x + 1.5)^2 + (y + 5)^2 = -70.5\]\[\implies -\frac{2(x + 1.5)^2}{70.5} - \frac{(y + 5)^2}{70.5} = 1\] @mukushla Is it just me or is this a degenerate conic?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yes it is...

OpenStudy (anonymous):

Alright. Thanks. Just wanted to make sure :)

OpenStudy (anonymous):

yw :)